OpenStudy (anonymous):

Find the point on the curve y= x sqrt (x) where the tangent line is parallel to the line 6x - y = 4

OpenStudy (anonymous):

i first found the slope of the line y= 6x -4 so the slope is 6 the derivative is y' = 3/2 x ^1/2 (using power rule) and so i did 3/2 x^1/2 = 6 but i couldnt get the correct answer, why?

OpenStudy (anonymous):

if it is \[x\sqrt{x}\] derivative is \[\frac{3}{2}\sqrt{x}\] then set \[\frac{3}{2}\sqrt{x}=6\] \[\sqrt{x}=6\times \frac{2}{3}=4\] \[x=16\]

OpenStudy (anonymous):

ohhhhh i caught my mistake!! thanks